Daily Reports

Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote mode.

Jan 12th, Tuesday

16h - 21h; Bilal, Angel: Started about 15h (UT); tau(225GHz)~0.15, stable conditions. One beammap (s161: tau~0.13; El.~41 deg.) & calib1scan on Uranus. Pointing & calib1scan on secondary calibrators (MWC349, CRL2688, NGC7027) and 2251+158. Target tmc1-cp (project 096-20) observed for 1.5h. Project finished.

21h - 03h; Angelos, Suzanne, Hélène: Calibrations done: focus twice and skydip once. Atmospheric conditions: tau between ~ 0.12 and 0.17, stable pointing corrections and beams (2mm FWHM ~ 18"), relatively small sky fluctuations. Observed sources (program 160-16): NGC925 (8 scans, ~ 177 min), NGC2976 (6 scans, ~ 49 min), and NGC5194 (2 scans, ~ 48 min). We lost about 1h15 at the beginning of the shift because of a difficult setup of all the needed windows. Initially, they were split up between VNC 30 and VNC 32, which is not practical when observing remotely with a tiny laptop screen. Many thanks for the help received in the process of setting everything up nicely on VNC 32! We also lost time when switching to the pointing quasar for our last source, and back, because of the azimuth limits. Will know better next time.

Jan 13th, Wednesday

03h - 09h; Longji: The observation started at 3:20 LT when the last scan of the previous shift was completed. I first observed COSMOS until 5:45 LT. After doing focus on 0851+202, I switched to GOODS-N until the end of the shift. The pointing was checked every 1 hour. The weather was clear and tau values were around 0.15 during the shift.

09h - 15h; Jean-François: Observed GASTON 122-16_dust sources (top priority source L1689b) for 3 hours and then GASTON 122-16_GP (cycling thru the 3 sources) for 1 hour. Excellent weather through the run. Only issue: focusing was difficult at 13h (LT). It took a source away from the sun to succeed. The telescope had been observing the whole morning at about 30 deg from the sun.

15h - 21h; Angel: Continuing on GP_L24p5 of GASTON, and then 1 rep. on GP_L23p9. Point and focus on Uranus + 2 skydips. After ~1.5h the antenna was relaxed enough to obtain beam sizes of 12.8" at 1.3 mm and 18.5" at 2 mm. Changing to project 183-20 at ~17h UT. A total of 10 rounds (~2.5h) gathered on MOO0319. Tau below 0.21 and quite a stable atmosphere during the shift.

21h - 03h; Angelos, Suzanne, Hélène: about calibrations, focus three times (-0.44, -0.13, and -0.40) and skydip once. Atmospheric conditions: tau between ~ 0.22 and 0.3, stable pointing corrections and beams (2mm FWHM ~ 18"), acceptable sky fluctuations. Observed sources (160-16): NGC925 (10 scans, ~ 148 min) and NGC2146 (14 scans, ~ 123 min). This time, everything went smoothly, but opacities rose steadily, with occasional little peaks. As far as we can tell, the atmosphere was stable enough. However, at some point, PIIC stopped reducing the 1mm data (but kept producing results at 2mm). We also had trouble finding suitable programs for opacities ~ 0.3 at the end of the shift. Note: for once, the taumeter measurements are not in line with the predictions. Usually, they are in very good agreement. We realized too late that a beammap was needed, i.e. when Uranus was getting low. We'll do it next night if possible, unless it's done before.

Jan 14th, Thursday

03h - 09h; Guilaine, Angel: Tau was slowly but surely increasing from 0.28 to 0.35. Target IC443G of project 134-20 observed for ~1.5h until reaching -almost simultaneously- the elevation and opacity limits (tau<0.3) provided by the PI. After this, there were no targets of scheduled NIKA2 Pool projects above the horizon and allowing opacity upper limits > 0.3. At 05h20m (LT) we switched to the EMIR backup project 171-20. Targets observed: 8942-12702, 9181-12704, 8247-6103, 8997-1902. Opacity approximating to 0.4 during the last 2 hours of the shift. This seems to be the trend for daylight hours.

09h - 15h; Jean-François: Continue EMIR from the last run with Angel's assistance (2 hours). Then 126-20 with external calibrator turned on and Bilal's assistance (2 hours). Then 122-16 (GP) (1 hour) with weather conditions being at the requested limit (Note: "external calibrator" was purposedly left turned on during 122-16 under the assumption it was not hurting the data and it was useful for calibrator signal implementation into analysis software).

15h - 21h; Angel: Tau was about 0.35 with several spikes during the shift. Continuing on GP_L24p5 and then GP_L23p9 targets of GASTON. A total of 10 maps (126-20) of the Galactic equator were obtained after that. The shift concluded with one repetition on IC443G (134-20).

21h - 03h; Angelos, Suzanne, Hélène for programs 134-20 and 160-16. Calibrations done: focus three times (-0.15, -0.34, and -0.52) and skydip twice, one beammap on Uranus with tau ~ 0.25. Atmospheric conditions: tau between ~ 0.3 and 0.4 with sharp peaks while on IC443G, then decreased quickly and remained stable around 0.23 - 0.25 for the rest of the shift. Observed sources: IC443G (project 134-20): 4 scans. Uranus beammap. NGC2146: 16 scans -140 min. One pointing scan (20210115s7) could not be reduced because of a corrupted imbfits file, and had to be repeated. Everything else went smoothly.

Jan 15th, Friday

03h - 09h; Nico for N2CLS. Tau was 0.2 or slightly above, with the forecast predicting improvement. Therefore, I took the risk to observe Goodsnorth. I chose Goodsnorth rather than Cosmos because they were at comparable elevation, but GN is smaller, hence less sensitive to atm variations, and it was rising in elevation. Opacity did decrease a bit before rising again and fluctuating. I did a focus at the beginning of my shift and controlled it on a pointing at about half the shift. Everything went smoothly.

09h - 15h; Jean-François: Excellent weather throughout the run (tau_225 ~0.1) as predicted by Bilal-implemented predictor. Aurore Backmann joined for the run. First, oversee Deepfield LP operation (40 minutes), then GASTON LP 122-16 (L1689b) (3hours), then GASTON LP 122-16 (GP) (2 hours).

15h - 21h; Angel (joined for the run Yu Cheng, Lingrui Ling, and Denis Burgarella as trainees): outstanding opacity (0.08 < tau < 0.11) and stability during the shift. Continuing on L23p3 and then on L24p5 (GASTON). The latter one observed twice by mistake, sorry. Observed sources: ACTJ0119, project 199-16, 5 repetitions (1h40m); L1506, project 094-20, 2 repetitions (40m); Sh2-289, project 118-20, 1 repetition (1h).

21h - 03h; Angelos, Suzanne, Hélène. Calibrations: calib_1scan on Uranus with tau ~ 0.1, focus three times. Atmospheric conditions: excellent ! with tau between 0.10 and 0.13. Observed sources: NGC925: 10 scans, NGC2146: 6 scans, NGC5194: 3 scans. We started at ~ 9:20 after the last scan of the previous project ended and had excellent observing conditions. But the frequency of corrupted files (twice during pointing, once during focus) seems to increase. We could not check on target scans because their processing is too long.

Jan 16th, Saturday

03h - 09h; Nico for N2CLS. After a focus/pointing on 1055+018, I've observed COSMOS for a bit less than 2 hours. Opacity was about 0.1 when I started, has been increasing ever since. I now leave the antenna to EMIR (~5h20). Frédéric Damour observed 9182-9102 of the backup project 171-20. Telescope stopped at 07h due to high wind speed.

09h - 15h; Andrew (on-hold): Telescope stopped. High wind speed.

15h - 21h; Karine, Isabelle: We have started observing at 19h30 (local time) with pointing and focus on 0316+413. Opacity ~0.25 - 0.28 (spike at 0.32); the beam value was a bit large at 1mm. We observed the source L1506 of project 094-20: 2.5 NReps (i.e ~ 50mn).

21h - 03h; Suzanne, Angelos, Hélène. Calibrations: beammap on Uranus with tau ~ 0.3 or slightly below, calib_1scan on Uranus, skydip, focus four times atmospheric conditions: tau around 0.3 in the first half of the shift, then between ~ 0.23 and 0.25, with occasional spikes; atmosphere reasonably stable as far as we can tell. Observed sources: NGC1331_1 for program 138-20: 4 scans, NGC925: 4 scans, NGC2146: 10 scans, NGC2841: 2 scans. We started with the telescope out of focus, so had to repeat the focus sequence to obtain a good fit. Note that we sometimes forgot to switch projects, so that two focus observations accidentally ended up being registered under 138-20 and 160-16, respectively. We wrote comments in TAPAS about this. At the end of the shift, we tried to point near our last source, NGC2841, at ~ 70 degrees elevation, but were unable to obtain decent pointing corrections. The beamwidth looked OK, but the two quasars we tried looked too faint (even with the "f" option in nkpoint). We chose to rely on the pointing model without entering the pointing corrections.

Jan 17th, Sunday

03h - 09h; Ankur, Kaustuv: The conditions were stable throughout the shift and tau opacity was below 0.25 overall. We observed 4 repetitions on project 180-20 on source eFEDS3M. Then moved to project 192-16 and observed 2 repetitions on source GOODSNORTH.

09h - 15h; Andrew: It took some time to focus the telescope and recover good beams at 1mm, but eventually we could begin at ~11h. It's not clear to me what the issue was, but observations for GASTON / 122-16 went smoothly after this. Tau was steady at the ~0.25 limit for this project. We observed L1689B for 4 scans, and took a total of 8 scans on the GP fields.

15h - 21h; Isabelle, Karine: We have started our shift with thermal stress of the telescope: the beam shapes and FWHM were bad, and it took about 3 hours to get back to the nominal values. We have performed observations for the time filler project (sources: GP072_0+0_0 ; GP076_3+0_0, with 10mn / source); On our project, 094-20, we completed the observations of G155 (total ~1h), and performed 3 Nreps for L1506 (total ~ 1h20).

21h - 03h; Suzanne, Golshan, Hélène: Calibrations: calib_1scan on Uranus, focus four times, skydip. Atmospheric conditions: tau fluctuating between ~0.26 and 0.32 during the first part of the shift, then more stable at ~0.25 - 0.27. Observed sources: NGC1333_1 (program 138-20): 12 scans, can be considered completed; NGC925 (160-16): 1 scan, probably not useable; NGC2841 (160-16): 10 scans, is probably completed. The opacity was too high in the first part of the shift for a beammap on Uranus. When we moved to N925, we launched two scans, but only the first was executed. The second one was loaded but not executed, and had to be cancelled. Then we moved to N2841 because N925 was getting too low in elevation. Since N2841 got over 70 degrees in elevation, we gradually decreased the scan speed from 60 to 40 arcsec/s.

Jan 18th, Monday

03h - 09h; Ankur, Kaustuv: We started with opacity around 0.26, with fairly stable conditions. We did 3 repetitions on project 180-20; opacity had some spikes and fluctuations up to 0.44. Stability reduced as well. Telescope was then handed over to Frédéric to use EMIR for backup program 171-20. At 08h (LT) the instrument was switched back to NIKA-2, and opacity continues to be above 0.25. Moved to project 126-20 for Galactic plane scans.

09h - 15h; Andrew: Tau stayed fairly stable between ~0.25 - 0.30 for the whole time. Unfortunately, this is just beyond the limit for GASTON, so project 126-20 was observed.

15h - 21h; Isabelle, Karine: We had issues with the beam shape and FWHM values and didn't get the expected nominal values. For this reason, and according to the mean opacity values (tau > 0.25), we observed the time filler project 126-20.

21h - 03h; Suzanne and Hélène: Calibrations: calib_1scan and beammap on Uranus, focus three times. Atmospheric conditions: tau fluctuating between 0.22 and 0.3 during the first part of the shift, then decreased slowly and got more stable until reaching 0.21 at the end of the shift; sky reasonably stable. Sources observed: for program 138-20, NGC2068: 4 scans; for program 160-16, NGC925: 4 scans, can be considered completed (?), NGC3198: 14 scans, and NGC5194: 1 scan. We did our third beammap since Jan. 12, and this one should be the best of the three, since it was taken at tau ~ 0.22 to 0.26, i.e. lower than the first two times, and the atmosphere was not unstable, as far as we can gauge from the PIIC windows. On NGC3198, we began at 60"/s and decreased the scan speed to 50 and then 45"/s, as it was getting higher in elevation. The lone scan on NGC5194 was done to complete the other lone scan taken a few nights ago, in the orthogonal scan direction. Then it was already 3am. We still had trouble with the RTA of the focus scans (in IDL). Once, the middle scan could not be processed, but we obtained nevertheless a good fit from the remaining 4 scans.

Jan 19th, Tuesday

[note by Stefano from Grenoble: the PIIC monitor was not processing the Ar1 and Ar3 data because the make-imbfits script for Ar1 and 3 stopped. We realized it yesterday. Angel has activated it again (thank you!!!). It will take time to catch up and the observers might notice that the monitor is processing older scans. If you notice that it stops again, please inform the AoD. Thank you]

03h - 09h; Ankur, Kaustuv: We began with a sky dip. Tau ~ 0.20 at the start, but wind speed and instabilities were gradually increasing. We started with project 180-20, but cancelled scans due to high wind speeds. After a break, and lower wind speed we moved to project 192-16 on GOODSNORTH source and did 2 repetitions on it. At the end of shift tau was 0.14 and stable atmospheric conditions overall.

09h - 10h; Angel: Tau~0.12, stable. Continuing on GOODSNORTH for a while. After sunshine, I started with 199-16: 2 repetitions on PSZ2G087.

DailyReportsNika2Pool120121 (last edited 2021-01-22 12:08:19 by NikaBolometer)